# Subtleties in the pseudomodes formalism

### v1

@@ IV Many-Mode Fitting
If we are willing to use a large number of pseudomodes, we can ensure that couplings to the residual baths remain small. However, it stops being practical to run optimization algorithms for fitting. In this case, we might proceed as in Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025). The procedure uses a diagonal configuration, so each pseudomode couples only to the system and to its own residual bath and the effective spectral density is given by (18). There are a few benefits to this many-mode procedure. It is more efficient to set up since no optimization routine is required, though computation with a large number of modes is expensive. Additionally, by using a large number of uncoupled modes at different energies, analyzing the currents between individual pseudomodes and the system can provide information about energy currents. For each bath, the procedure is as follows, where the index $\alpha$ is suppressed:
@@
Figure 4: Effective spectral density obtained by applying the many-mode approximation procedure described above to a semi-elliptical spectral density, for 100 pseudomodes. The effective spectral density oscillates rapidly about the true spectral density.
@@ IV.1 Proof of Nonconvergence
To show where the issue emerges, consider $J^{\rm eff}_{ij}(\varepsilon_{k})$, the effective spectral density at one of the pseudomode energies. First, note that we can write:
@@
where the $\gamma$’s conveniently cancel since $\varepsilon_{q}-\varepsilon_{k}=(q-k)\gamma_{\alpha}$. Now, assuming a very large number of pseudomodes, we can approximate $J^{\rm eff}_{ij}(\varepsilon_{q})$ in (59) by
@@
where the shorthand $J^{\prime}=dJ/d\omega$. We then can see that the contributions from the linear terms at $\varepsilon_{k+q}$ and $\varepsilon_{k-q}$ will cancel each other, so we can rewrite (59) as:
$J^{\rm eff}_{ij}(\varepsilon_{k})=\sum_{q}\frac{J_{ij}(\varepsilon_{k})}{2\pi\left[(q-k)^{2}+1/4\right]}\,.$ (61)
We can arrive at (61) more rigorously if we restrict the sum to a set number of terms ($q$ such that $|k-q|\leq 8$ is sufficient to show nonconvergence) and use the fact that $J_{ij}$ is a continuous function to show that we can always choose a number of modes so that $J_{ij}(\varepsilon_{k})-J_{ij}(\varepsilon_{q})<\epsilon$ for any $\epsilon$. In the limit of infinitely many pseudomodes we can take $\epsilon\to 0$ and the limits on the summation to $\pm\infty$ and likewise arrive at (61).
@@
Figure 5: Effective spectral density for pseudomodes in a diagonal configuration (blue) and for modes in 2x2 non-diagonalizable blocks (pink), compared to the error expected in the infinite-mode limit at the pseudomode energies (above $J_{\rm eff}$) and halfway between them (below $J_{\rm eff}$, for two example spectral densities. 100 pseudomodes are used in each figure, creating an effective spectral density that is either the sum of 100 Lorentzians or 50 squared Lorentzians. Note that where the effective spectral density differs from the infinite-mode prediction, it is where the second derivative of the true spectral density is large, so the first-order Taylor approximation (60) is locally a poor approximation. The left spectral density is semi-elliptical, and the right is a sum of two Lorentzians.
@@ IV.2 Many-Mode Approach with Non-diagonalizability
    Choose an even number of pseudomodes $n$ such that $\gamma=4\left(\frac{\omega_{M}-\omega_{m}}{d/2-1}\right)$ is an acceptably small coupling to the residual baths.
@@
We can numerically compare the two methods to see these differences in error, as shown in Fig. 5.
@@ V Connection to Scattering Theory
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p}\tau_{\alpha k,p}^{2}\delta(\omega-\omega_{p})\,$, (76)
@@
$\big(\omega-H_{S}-\zeta_{L}^{\dagger}\mathcal{F}_{L}\zeta_{L}-\zeta_{R}^{\dagger}\mathcal{F}_{L}\zeta_{R}\big)\mathcal{G}_{S}=1\,.$ (99)
@@
Next we express the four quantities $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$ [which appear in (81)-(84)] in terms of $\mathcal{G}_{S}$. Combining (92) with (97) we get
$\mathcal{G}_{LR}=\mathcal{F}_{L}\zeta_{L}\mathcal{G}_{S}\zeta_{R}^{\dagger}\mathcal{F}_{L}\,.$ (101)
Combining (93) with (95) we get
$\mathcal{G}_{RL}=\mathcal{F}_{L}\zeta_{R}\mathcal{G}_{S}\zeta_{L}^{\dagger}\mathcal{F}_{L}\,.$ (102)
@@
We therefore see that we can write this in a form similar to (73), provided we define $n_{S}\times n_{S}$ dimensional matrices
$\mathcal{J}_{\alpha k}^{+}=\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}^{\dagger}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}\zeta_{\alpha}\,,\qquad\mathcal{J}_{\alpha k}^{-}=\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}^{\dagger}\zeta_{\alpha}\,.$ (106)
Then
@@
As usual, we now take the pseudomodes’ residual spectral densities to be constant:
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha k}\,.$ (108)
Then, if we let $\mathcal{J}_{\alpha}^{+}(\omega)=\sum_{k}\mathcal{J}_{\alpha k}(\omega)$, we can see that, starting from (106):
$\mathcal{J}^{\pm}_{\alpha}$ $=\zeta_{\alpha}^{\dagger}\frac{1}{\omega+iW^{\dagger}_{\alpha}}\Gamma_{\alpha}\frac{1}{\omega-iW_{\alpha}}\zeta_{\alpha}$ (109)
@@
Likewise, if we have constant residual spectral densities (108) then $\Sigma_{\alpha}=i\Gamma_{\alpha}/2$}, and so the effective system Green’s function is:
@@
where $\mathcal{G}_{S}$ is just the NEGF that we would calculate from the system interacting with a set of baths with spectral densities $J^{\rm eff}_{\alpha}$. Thus, we can see that so long as the spectral densities (or correlation functions) match exactly, the transmission functions will also match exactly.
@@ Acknowledgments
This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. L.P.B. acknowledges Research Ireland for support through the Frontiers for the Future project. Optimization was performed with the Optim.jl Julia package Mogensen and Riseth (2018).
@@ Appendix A Derivation of the Effective Spectral Density
Appendix B Inversion Problem for 2 Pseudomodes
@@
In this appendix we explore the method presented in section III for the simple case of two pseudomodes. Suppose we obtain a 2-mode fit from Prony’s method with eigenvalues $\varepsilon_{1},\varepsilon_{2},\gamma_{1},\gamma_{2}$ and amplitudes $\kappa_{1},\kappa_{2}$. First note that to satisfy (25), the $\kappa$’s must be of the following form:
@@ Appendix C Non-Diagonalizable Configuration as the Limit of a Diagonalizable One
$W$ is diagonalizable so long as $\epsilon\neq 0$, and its eigenvalues are
$\delta\left(1\pm i\sqrt{2\epsilon(1+\epsilon)}\right)\,.$ (159)
We then compute the memory kernel using (14) and obtain:

### v2

@@ IV Many-Mode Fitting
If we are willing to use a large number of pseudomodes, we can ensure that couplings to the residual baths are arbitrarily small. However, it stops being practical to run optimization algorithms for fitting. In this case, we might proceed as in Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025). The procedure uses a diagonal configuration, so each pseudomode couples only to the system and to its own residual bath and the effective spectral density is given by (19). There are a few benefits to this many-mode procedure. It is more efficient to set up since no optimization routine is required, though computation with a large number of modes is expensive. Additionally, by using a large number of uncoupled modes at different energies, analyzing the currents between individual pseudomodes and the system can provide information about energy currents. For each bath, the procedure is as follows, where the index $\alpha$ is suppressed:
@@
Note that the problem of non-convergence presented here arises from this particular method of choosing the pseudomode parameters. In general, one can always choose parameters to approximate a spectral density arbitrarily well, even in the diagonal (uncoupled) case Trivedi et al. (2021).
Figure 5: Effective spectral density obtained by applying the many-mode approximation procedure described above to a semi-elliptical spectral density, for 100 pseudomodes with $\omega_{m/M}=\pm 1.4$. The effective spectral density oscillates rapidly about the true spectral density.
@@ IV.1 Proof of Nonconvergence
To show where the issue emerges, consider $J^{\rm eff}_{ij}(\varepsilon_{k})$, the effective spectral density at one of the pseudomode energies. Since $\gamma$ is defined as the spacing between pseudomode energies $(\varepsilon_{k}-\varepsilon_{k-1})$, we can write (59) evaluated at $\varepsilon_{k}$ as:
@@
where the $\gamma$’s conveniently cancel since $\varepsilon_{q}-\varepsilon_{k}=(q-k)\gamma$. Now, assuming a very large number of pseudomodes, we can approximate $J^{\rm eff}_{ij}(\varepsilon_{q})$ in (60) by
@@
where the shorthand $J^{\prime}=dJ/d\omega$. This approximation is reasonable in the case that we have a large number of pseudomodes, since terms in the sum where |qk||q-k| is large are suppressed quadratically. We then can see that the contributions from the linear terms at $\varepsilon_{k+q}$ and $\varepsilon_{k-q}$ will cancel each other, so we can rewrite (60) as:
$J^{\rm eff}_{ij}(\varepsilon_{k})=\sum_{q=1}^{n}\frac{J_{ij}(\varepsilon_{k})}{2\pi\left[(q-k)^{2}+1/4\right]}\,.$ (62)
We can arrive at (62) more rigorously if we restrict the sum to a set number of terms ($q$ such that $|k-q|\leq 8$ is sufficient to show nonconvergence) and use the fact that $J_{ij}$ is a continuous function to show that we can always choose a number of modes so that $J_{ij}(\varepsilon_{k})-J_{ij}(\varepsilon_{q})<\epsilon$ for any $\epsilon$. In the limit of infinitely many pseudomodes we can take $\epsilon\to 0$ and the limits on the summation to $\pm\infty$ and likewise arrive at (62). This scheme works even at points where the linear approximation (61) is poor, since it only relies on the spectral density being continuous.
@@
Figure 6: Effective spectral density for pseudomodes in a diagonal configuration (blue) and for modes in 2x2 non-diagonalizable blocks (pink), compared to the error expected in the infinite-mode limit at the pseudomode energies (above $J_{\rm eff}$) and halfway between them (below $J_{\rm eff}$, for two example spectral densities. 100 pseudomodes are used in each fit in each figure, creating an effective spectral density that is either the sum of 100 Lorentzians (100 uncoupled modes) or 50 squared Lorentzians (100 modes, coupled as 50 pairs), with $\omega_{m,M}=\pm 1.4$ for both figures. Note that where the effective spectral density differs from the infinite-mode prediction, it is where the second derivative of the true spectral density is large, so the first-order Taylor approximation (61) is locally a poor approximation. The left spectral density is semi-elliptical, and the right is a sum of two Lorentzians.
@@ IV.2 Many-Mode Approach with Non-diagonalizability
    Choose an even number of pseudomodes $n$ such that $\gamma=4\left(\frac{\omega_{M}-\omega_{m}}{d/2-1}\right)$ is an acceptably small coupling to the residual baths (with the factor of $4$ included because the squared Lorentzians have width $\delta=\gamma/4$).
@@
We can numerically compare the two methods to see these differences in error, as shown in Fig. 6. This oscillating error can be further mitigated by increasing $\gamma$ and adjusting the elements of $\zeta$ accordingly. However, while this reduces the magnitude of the oscillations, it makes the approximation (61) (and the corresponding approximation in the two-layer case) worse, and fails to fit the spectral density anywhere its second derivative is too large. It is likely that this issue could be mitigated by further refining this many-mode fitting procedure.
@@ V Connection to Scattering Theory
The method of pseudomodes is typically used to treat interacting systems. However, in this section we turn our attention to the methods of non-equillibrium Green’s functions (NEGF) and scattering theory, a common method for non-interacting systems Blanter and Büttiker (2000); Büttiker (1992); Levitov et al. (1996); Esposito et al. (2009), to show how the same pseudomode structure and effective spectral density that we derived using quantum master equations emerges in the context of scattering theory as well. The results of this section are largely independent of the previous section. They have been included here for two reasons. First, they illustrate how the same pseudomode structure can emerge from an entirely different approach. Second, they allow one to compute transmission functions resolved in the different lead modes. This makes it possible to study the tunneling amplitudes between the two baths, through the system, but resolved in the different lead modes, possibly providing new insights into the problem.
@@
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p\in(\alpha,k)}\tau_{\alpha k,p}^{2}\delta(\omega-\omega_{p})\,$, (77)
@@
$\left(\omega-H_{S}-\zeta_{L}^{\dagger}\mathcal{F}_{L}\zeta_{L}-\zeta_{R}^{\dagger}\mathcal{F}_{L}\zeta_{R}\right)\mathcal{G}_{S}=1\,.$ (100)
@@
Next we express the four quantities $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$ [which appear in (82)-(85)] in terms of $\mathcal{G}_{S}$. Combining (93) with (98) and (94) with (96) we get
$\mathcal{G}_{LR}=\mathcal{F}_{L}\zeta_{L}\mathcal{G}_{S}\zeta_{R}^{\dagger}\mathcal{F}_{L}\,,\quad\mathcal{G}_{RL}=\mathcal{F}_{L}\zeta_{R}\mathcal{G}_{S}\zeta_{L}^{\dagger}\mathcal{F}_{L}\,.$ (102)
@@
We therefore see that we can write this in a form similar to (74), provided we define $n_{S}\times n_{S}$ matrices
$\mathcal{J}_{\alpha k}^{+}=-\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}^{\dagger}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}\zeta_{\alpha}\,,\qquad\mathcal{J}_{\alpha k}^{-}=-\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}^{\dagger}\zeta_{\alpha}\,.$ (106)
Then, finally:
@@
As in (14), we now take the pseudomodes’ residual spectral densities to be constant:
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha,kk}\,.$ (108)
Then, if we let $\mathcal{J}_{\alpha}^{+}(\omega)=\sum_{k}\mathcal{J}_{\alpha k}(\omega)$, we can see that, starting from (106) with $W_{\alpha}$ as in (16):
$\mathcal{J}^{\pm}_{\alpha}(\omega)$ $=-\zeta_{\alpha}^{\dagger}\frac{1}{\omega+iW^{\dagger}_{\alpha}}\Gamma_{\alpha}\frac{1}{\omega-iW_{\alpha}}\zeta_{\alpha}$ (109)
@@
Likewise, if we have constant residual spectral densities (108), then $\Sigma_{\alpha}=i\Gamma_{\alpha}/2$}, and so the effective system Green’s function is:
@@
where $\mathcal{G}_{S}$ is just the NEGF that we would calculate from the system interacting with a set of baths with spectral densities $J^{\rm eff}_{\alpha}$. Thus, we can see that so long as the true and effective spectral densities (or correlation functions) match exactly, the transmission functions (74) and (118) will also match exactly.
@@ Acknowledgments
This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. L.P.B. acknowledges Research Ireland for support through the Frontiers for the Future project. Optimization was performed with the Optim.jl Julia package Mogensen and Riseth (2018). No generative AI tools were used in the writing of this article.
@@ Appendix A Derivation of the Effective Spectral Density
Appendix B Memory Kernel Fitting Methods
Suppose we have a memory kernel $\chi(t)$ that we want to approximate as:
$\chi(t)=\sum_{k=1}^{n}\kappa_{k}e^{-\lambda_{k}t}\,$, (133)
for complex parameters $\kappa_{k},\lambda_{k}$, where we want $n$ as small as possible to get a good fit. In this section we describe two algorithms from discrete signal processing, Prony’s method and ESPRIT, that have been used for this fitting problem.
B.1 Prony’s Method
Prony’s method de Prony (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) consists of the following steps:
  1. 1.
    Choose a time interval $\Delta t$ and cutoff time $t_{c}=2M\Delta t$ to sample the correlation function, where $M\geq 2n+1$.
  2. 2.
    Let $\chi_{j}=\chi(j\Delta t)$ for $j=0,1,\dots,2M$ and construct the $(M+1)\times(M+1)$ Hankel matrix $\mathbf{H}_{ij}=\chi_{i+j+1}$.
  3. 3.
    Compute the Takagi decomposition $H=U\sigma U^{T}$, with $\sigma$ diagonal and $U$ unitary Ikramov (2012), noting that we take the transpose $U^{T}$ rather than the adjoint $U^{\dagger}$. This yields a set of normalized vectors $\mathbf{u}_{m}$ such that $\mathbf{H}\mathbf{u}_{m}=\sigma_{m}\mathbf{u}_{m}^{*}$. This is analogous to finding the eigen-decomposition of a Hermitian matrix as $\mathbf{H}{|{m}\rangle}=\sigma_{m}{|{m}\rangle}$, but for a complex-symmetric matrix. Order the $\sigma_{m}$ by descending values of $|\sigma_{m}|$ and the $\mathbf{u}_{m}$ accordingly.
  4. 4.
    Define the polynomial $F(z)=\sum_{j=0}^{M}\left(\mathbf{u}_{n}\right)_{j}z^{j}$ and obtain its roots $f_{0},f_{1},\dots f_{N}$ ordered by increasing modulus.
  5. 5.
    Of these roots, the first $L$ will have $|w_{k}|<1$. Write these roots as $r_{k}e^{i\theta_{k}}$ with $\pi<\theta\leq\pi$ and choose the exponents in (133) to be
    $\lambda_{k}=-\frac{2M}{t_{c}}\left[\log(r_{k})+i\theta_{k}\right]\,.$ (134)
  6. 6.
    Obtain the amplitudes $\kappa_{k}$ by least-squares fitting the equations
    ${\chi_{j}=\sum_{k=1}^{n}\kappa_{k}w_{k}^{j}}\,,\quad j=0,1,\dots n$ (135)
B.2 ESPRIT Algorithm
The ESPRIT (Estimation of Signal Parameters by Rotational Invariance Techniques) algorithm Paulraj et al. (1986); Roy et al. (1986); Sahnoun et al. (2017) is designed to be less numerically unstable and more robust to noise than Prony’s method. It consists of the following steps:
  1. 1.
    Choose a time interval $\Delta t$ and cutoff time $t_{c}=M\Delta t$ to sample the correlation function, where $M\geq n+1$.
  2. 2.
    Let $\chi_{j}=\chi\big((j-1)\Delta t\big)$ for $j=1,\dots,2M+1$.
  3. 3.
    Choose the parameter $n+1\leq L<M$, let K=ML+1K=M-L+1, and construct the $L\times K$ Hankel matrix $\mathbf{H}_{ij}=\chi_{i+j-1}$.
  4. 4.
    Compute the singular value decomposition $H=U\Sigma V$, where $U$ is the $L\times L$ matrix of left singular vectors, $V$ is $K\times K$, and $\Sigma$ is the $L\times K$ matrix of singular values.
  5. 5.
    Define $U_{s}$ to be the $L\times n$ matrix obtained by taking the $n$ columns of $U$ corresponding to the $n$ largest singular values, and construct the matrix $F$ by solving
    $\underline{U_{s}}F=\overline{U_{s}}\,$, (136)
    where $\underline{U_{s}}$ ($\overline{U_{s}}$) is the $(L-1)\times n$ matrix obtained by taking the matrix $U_{s}$ and removing the bottom (top) row. (Note that while $U$ is unitary, the matrices $U_{s},\overline{U_{s}},\underline{U_{s}}$ are not unitary).
  6. 6.
    Compute the eigenvalues of $F$, denoted $f_{k}$. These relate to the exponents in (133) by:
    $f_{k}=e^{-\lambda_{k}\Delta t}\,.$ (137)
  7. 7.
    Obtain the amplitudes $\kappa_{k}$ by least-squares fitting the equations
    $\chi_{j}=\sum_{k=1}^{n}\kappa_{k}f_{k}^{j}\,,\quad j=0,1,\dots n\,.$ (138)
Fitting with either of these methods can be sensitive to the choice of discretization made in step 1, though ESPRIT is far more stable than Prony’s method and can handle a larger number of samples of $\chi(t)$. Thus, it is often practical to sample a large number of interval choices and then select the choice that gives the effective spectral density that best fits the true spectral density, or to perform optimization on the endpoints of the time interval used.
$W$ is diagonalizable so long as $\epsilon\neq 0$, and its eigenvalues are
$\delta\left(1\pm i\sqrt{2\epsilon(1+\epsilon)}\right)\,.$ (140)
We then compute the memory kernel using (15) and obtain:
[…]
$\mathfrak{Re}(\kappa_{1})$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}+\frac{2\zeta_{1}\zeta_{2}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (144)
$\mathfrak{Re}(\kappa_{2})$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}-\frac{2\zeta_{2}\zeta_{1}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (145)
$\mathfrak{Im}(\kappa_{1})$ $=\frac{\zeta_{1}^{2}-\zeta_{2}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,$, (146)
$\mathfrak{Im}(\kappa_{2})$ $=\frac{\zeta_{2}^{2}-\zeta_{1}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,.$ (147)
Note that each of these diverges as $\epsilon\to 0$. However, if we construct the spectral density from (27), we obtain:
Taking the limit of $\epsilon\to 0$, we obtain:
exactly as before.
Appendix D 3x3 non-diagonalizable Configuration
We compute the eigenvalues of this matrix as:
$x=\sqrt{\frac{1}{2}\left(5\sqrt{5}-11\right)}\,$, (154)
and so we find a candidate 3x3 non-diagonalizable $W$ as:
and effective spectral density
for this pseudomode configuration.
Appendix E Inversion Problem for 2 Pseudomodes
In this appendix we explore the simple case of two pseudomodes. While a solution for the pseudomode parameters in this case has been previously shown Pleasance et al. (2020), we are able to show using the methods of section III that, for the 2-mode case, it is always possible to choose pseudomode parameters such that $\Gamma>0$ for effective spectral density that is nonnegative.
@@ Appendix E Inversion Problem for 2 Pseudomodes
Suppose we obtain a 2-mode fit to the memory kernel with eigenvalues $\varepsilon_{1},\varepsilon_{2},\gamma_{1},\gamma_{2}$ and amplitudes $\kappa_{1},\kappa_{2}$. First note that to satisfy (26), the $\kappa$’s must be of the following form:
@@
Appendix F Fit Parameters in Figures 2 and 4
Table 2: Fit parameters for the diagonal fit used in Fig. 2(a,b), as they appear in (20).
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.04934 0.47611 0.33539
0.04934 -0.47611 0.33539
0.06613 -0.16441 0.39853
0.06613 0.16441 0.39853
0.02641 0.7457 0.23523
0.02641 -0.7457 0.23523
Table 3: Elements $\Lambda_{ij}$ for the pseudomode configuration shown in Fig. 4, as they appear in (17).
$\Gamma$ $\zeta$ $\Lambda_{ij}$
-0.086 -0.222 - 0.318ii 0.0738 0.755-0.00389ii -0.131-0.00188ii 0.0696-0.122ii 0.164+0.237ii 0.0427-0.00337ii